News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 29 October 2004 9.00am EDT

During the last week Chandra completed the observing schedule as planned.

The IEPHIN instrument exhibited anomalous behavior on Oct 28 during a period of elevated operating temperature. The power supply voltage dropped as it became current limited due to the high temperature and is expected to return to its nominal value once the unit cools following a maneuver to a new attitude. This behavior has been seen before and mission planning constraints have been in use to avoid high temperatures. As a result, the constraints and temperature prediction software will be reviewed.

Of note last week was the Chandra Calibration Workshop and review held Oct 26-27. The workshop provided the community with the state of the art of the Chandra team's knowledge of the detectors and mirrors, and also encouraged the general community to participate in the process of calibrating the satellite (

A Chandra press release was issued on Oct 26 describing the detection of an extensive envelope of dark matter around an isolated elliptical galaxy. This discovery conflicts with optical data that suggest a dearth of dark matter around similar galaxies, and raises questions about how galaxies acquire and keep such dark matter halos. For details see:

The schedule of targets for the next week is shown below and includes an observation of M31 coordinated with HST.

M31-bulge              ACIS-I       Oct 30
AX J0207.7+3511        ACIS-S
2MASS 015727+13322     ACIS-S       Oct 31
PSR J1119-6127         ACIS-S
Radiation Belts
M101                   ACIS-S       Nov 1
F1255-3653             ACIS-S       Nov 2
F1300-2556             ACIS-S
PSR J1119-6127         ACIS-S
Orion Nebula Cluster   ACIS-S/HETG
Radiation Belts                     Nov 3
LBQS 0109-0128         ACIS-S       Nov 4
RXJ1856.5-3754         HRC-I
Sculptor Dwarf         ACIS-S
cl0230+1836            ACIS-I       Nov 5
SN2004et               ACIS-S
Radiation Belts                     Nov 6
Vega                   HRC-I
Vega (3 obs)           HRC-S
M101                   ACIS-S       Nov 7


All spacecraft subsystems continued to support nominal operations.

Return to Status Reports